Use activities to calculate the electrode potential of a 

hydrogen electrode in which the electrolyte is 0.0100 

M HCl and the activity of H2 is 1.00 atm.


1
Expert's answer
2021-03-30T07:49:02-0400

Ecell= K + 0.05916/0.0100 log (H+)

Concentration of HCl = 0.0100 M

Activity coefficient of H+ = 0.913

Pressure of H2 = 1.00 atm

Ecell = 0.913 + (0.05916/0.0100) log (H+) 11.2

= 0.93 + (5.916. ×-0.0395) 11.2

= 0.913 - 2.617 =

-1.254 V